# Scalable LinUCB: Low-Rank Design Matrix Updates for Recommenders with Large Action Spaces

### v1

@@ 4.2 Batch updates
Now we provide a generalization of Theorem˜1 to the case of batch updates. In this setting parameter is updated after some number of interactions is accumulated in the system. Formally, we write $X_{t}\in\mathbb{R}^{d\times B}$ for the batch of $B\in\mathbb{N}$ concatenated contexts collected during the $t$-th interaction round with the hand $a\in\mathcal{A}$. Note that, generally speaking, $B$ depends on $a$ and $t$, but we prefer to write $B$ instead of $B_{t,a}$ for notation simplicity.
@@
where the Cholesky-like factor $L_{t+1}$ is obtained recursively as
$L_{t+1}=L_{t}(I+Q_{t+1}Y_{t+1}Q_{t+1}^{\top}),$ (30)
with $Q_{t+1}\in\mathbb{R}^{d\times B}$ being a matrix with orthonormal columns and $Y_{t+1}\in\mathbb{R}^{B\times B}$ obtained using the fast symmetric factorization approach Ambikasaran described in Appendix˜B (see Theorem˜3).
@@
Proof of Theorem˜2 is provided in Section˜B.2. The shape $d_{t}$ of the matrices $U_{t}$ and $V_{t}$ can be inferred from the representations (31) - (2).
Note that the matrices $U_{t}$ and $V_{t}$ in (2) expand in the number of columns with time $t$, leading to excessive memory demands and significant computational overhead. At the same time, in practice, the context distributions associated with each action typically have low-rank structure. Below we show, how one can maintain the low-rank structure of $U_{t}$ and $V_{t}$, using the Projector-Splitting Integrator lubich2014projector .
@@ 5 Projector-Splitting Integrator
Now we aim to improve the LinUCB algorithm based on the representation (31)-(2). As already mentioned, the matrices $U_{t}$ and $V_{t}$ here grow with time $t$, which yields both computational and memory issues. At the same time, the key element of the representation (31) is the time-varying factor $D_{t}=U_{t}V_{t}^{\top}\in\mathbb{R}^{d\times d}$. In this case, we essentially observe the matrix dynamics $\{D_{t}\}_{t\in\mathbb{N}}$. Moreover, the matrices $D_{t}$ might be well-approximated by low-rank factors. One of the options for dynamic approximation of $D_{t}$ comes from a singular value decomposition
@@
Here the best rank-$r$ approximation is given by the truncated SVD. At the same time, the orthogonal factors $(\bar{U}_{t},\bar{V}_{t})$ and $(\bar{U}_{t+1},\bar{V}_{t+1})$, corresponding to the matrices $D_{t}$ and $D_{t+1}$, are not guaranteed to be close, which might yield additional computational instability. Instead, we propose in this paper to rely on the projector-splitting integrator (PSI) approach of lubich2014projector , which aims at a dynamical low-rank approximation objective
@@
where $\dot{D}_{t}=\frac{\mathrm{d}}{\mathrm{d}t}D_{t}$. The PSI approach enables efficient updates to the approximations $\tilde{D}_{t}$ in an iterative manner, without explicitly forming the matrices $\tilde{D}_{t}$. This algorithm maintains dynamical form of updates
@@
where the factors $\tilde{U}_{t},\tilde{V}_{t}\in\mathbb{R}^{d\times r}$ have orthonormal columns, and $S_{t}\in\mathbb{R}^{r\times r}$ is invertible. The matrix $S_{t}$ here can be diagonal, yet it is not required. Then, with computational complexity of QR decompositions of $\mathbb{R}^{d\times r}$ matrices and multiplication of matrices of appropriate shapes, we obtain the factors $(\tilde{U}_{t+1},S_{t+1},\tilde{V}_{t+1}^{\top})$, such that
@@
We provide the implementation details in Algorithm˜4.
Now we explain the main steps of integrating Algorithm˜4 into the LinUCB algorithm with batch updates. At each batch update step, two matrices, $U_{t}$ and $V_{t}$, are maintained to represent the low-rank structure. Initially, these matrices are incrementally expanded by (27) or (2) as new data arrive, until their number of columns exceed a predefined rank threshold $r$. Once this occurs at some time step $t_{0}$, we compute the singular value decomposition (SVD) with fixed rank $r$ of the current product $U_{t_{0}}V_{t_{0}}^{\top}=\tilde{U}_{t_{0}}S_{t_{0}}\tilde{V}_{t_{0}}^{\top}$ to obtain the new factors $U_{t_{0}}=\tilde{U}_{t_{0}}S_{t_{0}}$, and $V_{t_{0}}=\tilde{V}_{t_{0}}$ for initialization Algorithm˜4. In subsequent batch updates, the matrices $U_{t}$ and $V_{t}$ are efficiently updated using the Algorithm˜4, which maintains the r-rank approximation. The corresponding increments $\Delta D_{t+1}$ defined by the algorithm’s update mechanism $U_{t+1}V_{t+1}^{\top}=U_{t}V_{t}^{\top}+\Delta D_{t+1}$ are directly obtained from (10) for the rank-1 update case:
@@
$\Delta D_{t+1}=Q_{t+1}(Y_{t+1}^{-1}+I)^{-1}Q_{t+1}^{\top}(I-U_{t}V_{t}^{\top}).$ (37)
The complete integration of this procedure within the LinUCB algorithm is summarized in Algorithm˜2. We highlight that we do not need to form the matrices $L_{t}^{-1}$ given in (31) explicitly.
Complexity analysis. Training stage. We analyze the computational complexity of the proposed PSI-LinUCB algorithm and we start from its training phase (described in Algorithm˜2). The computational complexity of PSI step
@@
does not depend on the number of interactions $B=B_{t,a}$ (that is, interactions with arm $a$ inside batch number $t$) and scales as $\mathcal{O}(dr^{2})$. This factor comes from the QR decomposition applied to $\mathbb{R}^{d\times r}$ matrices. Note that the complexity of computing SVD (line $11$ in Algorithm˜3) is also $\mathcal{O}(d(r+B)^{2}+(r+B)^{3})$. This can be achieved by computing the QR decomposition of the factors $U_{t},V_{t}$ of size at most $\mathbb{R}^{d\times(r+B)}$. The second dimension up to $r+B$ is due to line 10 in Algorithm˜3. Moreover, this SVD needs to be computed only once per each arm $a\in\mathcal{A}$. The last step contributing to asymptotic complexity is line $5$ in Algorithm˜3 with the computational complexity $\mathcal{O}(dB^{2}+B^{3})$ due to complexity of the QR and Cholesky decompositions. Thus, the overall complexity of handling blocks in the Update_arm algorithm (Algorithm˜3) is
@@
$\begin{split}a_{t}=\arg\max_{a\in\mathcal{A}_{t}}\left(\theta_{a}^{\top}x_{u,a}+\alpha\|(I-U_{t}V_{t}^{\top})x_{u,a}\|\right)\;.\end{split}$
@@
The primary computational cost of our algorithm arises during the training phase, which requires on average $\mathcal{O}(d(r^{2}/B+B)+r^{3}/B+B^{2})$ operations per interaction. By choosing $B\propto r$ and assuming $r\ll d$, this reduces to average complexity of $\mathcal{O}(dr)$ operations per iteration. To our knowledge, the $\mathcal{O}(dr)$ average complexity is the best computational cost achieved by LinUCB-based algorithms. This complexity is achieved by the algorithms suggested in the previous papers chen2020efficient , wen2024matrix .
@@
5:   $X_{t,a}.append\,[x_{u,a}]$
6:   $R_{t,a}.append\,[r]$
7:  end for
8:  for each arm $a$ in $batch_{t}$ do
9:   $U_{t+1,a},V_{t+1,a},\theta_{t+1,a}\leftarrow$ Update_arm($a$)
@@
11:end for
12:return $\theta_{n,a}$, $U_{n,a},V_{n,a}$ for each arm $a\in\mathcal{A}$.
@@ 6 Experimental Setup
Our experiments are conducted on four real-world datasets from Amazon product recommendations and classic Movielens 1M. In Sections 7.1 and 7.2, we demonstrate the efficiency of our algorithm compared to the batch LinUCB and LinUCB Classic for large context and arm spaces. In Section 7.3, we show that our algorithm can match the baseline performance without storing full-rank matrices, using only a small rank instead. In Section 7.4, we demonstrate its quality advantage over a CBSBFD algorithm with similar computational complexity.
Our results are conducted on $4$ real-world datasets from Amazon collection and the classical Movielens 1M dataset. Information about datasets is summarized in Table˜1.
All experiments are conducted on an Intel(R) Xeon(R) Gold 6348 CPU, 2.60GHz, kernel: 5.4.0-96-generic.
6.1 Baseline selection
We evaluate the performance of our proposed PSI-LinUCB algorithm against three baselines: the batch LinUCB, LinUCB with rank-1 updates via the Sherman–Morrison formula (LinUCB Classic), and the CBSCFD method proposed in chen2020efficient (see a detailed description of the CBSCFD method in Section˜3). Other approaches have several drawbacks, and we discuss below the reasons why we did not choose these algorithms as baselines.
The Frequent Directions (FD) sketching method kuzborskij2019efficient violates the positive-definite monotonicity property of the design matrices $A_{a}$, which may affect the algorithm’s stability. Therefore, we compare our approach against CBSCFD chen2020efficient , which addresses these limitations while maintaining the same $\mathcal{O}(md)$ computational and memory complexity. Similarly, chen2020efficient shows that CBSCFD outperforms both the CBRAP (Contextual Bandits via Random Projection) algorithm of yu2017cbrap and the Frequent Directions method of kuzborskij2019efficient . For this reason, we compare directly only with the stronger baseline method.
Dyadic Block Sketching wen2024matrix is an adaptive sketching technique whose theoretical advantages rely on restrictive assumptions about the spectrum of the design matrix, which may not hold in real-world data. Furthermore, in practice, adaptive sketching methods require choosing the initial sketch size $l_{0}$ and the error tolerance $\epsilon$ (see Algorithm in wen2024matrix ), and the performance of the algorithm is reported to be very sensitive to these hyperparameters. Specifically, under unfavorable spectral conditions, the sketch may grow excessively, erasing computational gains, whereas small values of $\epsilon$ degrade accuracy. This makes the approach operationally less stable than CBSCFD chen2020efficient . Moreover, the authors of wen2024matrix did not apply their method to real-world recommender system benchmarks, focusing only on classification problems.
Finally, Diag-LinUCB yi2023online is applicable only to contexts generated by a particular embedding model and cannot be applied to arbitrary feature representations. The method also reduces the context dimensionality to control latency, while our experiments focus on scalability with respect to high-dimensional contextual features.
@@ 6.2 Training Protocol
We divide the training process into phases:
  1. 1.
@@
6.3 Feature Extraction
User and item contexts are extracted using SVD to obtain. Given an interaction matrix $A\in\mathbb{R}^{m\times n}$ with $m$ users and $n$ items, we apply truncated SVD with rank $r^{\prime}=\min(r,\min(m,n)-1)$ to decompose it as $A\approx U\Sigma V^{T}$.
Item features are directly obtained from the right singular vectors after sorting by decreasing singular values: $V^{\top}\in\mathbb{R}^{n\times r}$. User features are then computed through matrix multiplication: $A\cdot V^{\top}\in\mathbb{R}^{m\times r}$.
SVD features are extracted from the data available during the warm-up period, which corresponds to the initial training split. To guarantee that each user and item in the test set has a valid context representation derived from this phase, all cold users are excluded from evaluation. This restriction arises because SVD-based feature extraction requires an existing interaction history. The use of SVD features enables the construction of high-dimensional contextual embeddings required for the experiments, however, our algorithm do not require the specific context representation and is capable of operating with arbitrary context.
6.4 Context Representation
The LinUCB algorithm is implemented with a batch processing approach for scalable recommendation systems. The algorithm maintains hybrid feature vectors ${x}_{ua}$ by taking the outer product of the user features $x_{u}\in\mathbb{R}^{d_{u}}$ and the item features ${x}_{a}\in\mathbb{R}^{d_{a}}$, resulting in $d=(r^{\prime})^{2}=\mathbb{R}^{d_{u}\times d_{a}}$ dimensional contexts.
7 Results
7.1 Scalability with context size
This experiment evaluates how the algorithms scale with the increasing context $d$ from 100 to 12,000.
We use the Amazon Magazine Subscriptions dataset, as it combines realistic user–item interactions with a moderate number of arms and measure computational time (training and evaluation) and memory usage while varying the SVD rank $r^{\prime}$ on a logarithmic scale. During the evaluation phase, we average the time and memory usage across months, as each month contains approximately the same number of interactions (see Section˜6.2), making it reasonable to consider the average time per month.
As shown in Figure˜1, the LinUCB requires $\mathcal{O}(d^{3})$ time for the warm-up phase, $\mathcal{O}(d^{2})$ time for inference, and $\mathcal{O}(d^{2})$ memory, since it maintains the full matrix $A_{a,t}$ for each arm $a$.
LinUCB Classic has both time and memory complexity of $\mathcal{O}(d^{2})$. The lower time complexity, compared to LinUCB, is achieved through rank-1 updates using the Sherman–Morrison–Woodbury formula.
In contrast, PSI achieves complexity $\mathcal{O}(dr^{2})$ time for the warm-up (see Remark˜1), $\mathcal{O}(dr)$ time for inference, and $\mathcal{O}(dr)$ memory. This reduction is possible because PSI does not store the full matrix $A_{a,t}$, but only its dynamically updated low-rank factors.

### v2

@@ 2.2 Batch updates
Now we provide a generalization of Theorem˜1 to the case of batch updates. In this setting parameter is updated after some number of interactions is accumulated in the system. Formally, we write $X_{t}\in\mathbb{R}^{d\times B}$ for the batch of $B\in\mathbb{N}$ concatenated contexts collected during the $t$-th interaction round with the arm $a\in\mathcal{A}$. Note that, generally speaking, $B$ depends on $a$ and $t$, but we prefer to write $B$ instead of $B_{t,a}$ for notation simplicity.
@@
where the Cholesky-like factor $L_{t+1}$ is obtained recursively:
$L_{t+1}=L_{t}(I+Q_{t+1}(M_{t+1}-I)Q_{t+1}^{\top}),$ (20)
with $Q_{t+1}\in\mathbb{R}^{d\times B}$ being a matrix with orthonormal columns and $M_{t+1}\in\mathbb{R}^{B\times B}$ obtained using the fast symmetric factorization approach Ambikasaran described in Appendix˜D (see Theorem˜3).
@@
Proof of Theorem˜2 is provided in Section˜D.3. The shape $d_{t}$ of the matrices $U_{t}$ and $V_{t}$ can be inferred from the representations (21) - (2).
2.3 Low-rank correction for $A_{t}^{-1}$
Note that the matrices $U_{t}$ and $V_{t}$ in (2) expand in the number of columns with time $t$, leading to excessive memory demands and computational overhead. We now motivate why the correction term for $\varepsilon^{-1}I-A_{t}^{-1}$ can be well approximated by a low-rank matrix. Consider the empirical covariance matrix $\sum_{s=1}^{t}x_{s}x_{s}^{\top},$ and let its SVD be given by
$\sum_{s=1}^{t}x_{s}x_{s}^{\top}=E_{x}\Sigma_{x}E_{x}^{\top}.$ (22)
Then the regularized design matrix and its inverse admit the representations
$A_{t}=E_{x}(\varepsilon I+\Sigma_{x})E_{x}^{\top},\qquad A_{t}^{-1}=E_{x}(\varepsilon I+\Sigma_{x})^{-1}E_{x}^{\top}.$ (23)
Then the correction term $R_{t}:=\varepsilon^{-1}I-A_{t}^{-1}$ writes as
$R_{t}=E_{x}\Bigl(\varepsilon^{-1}I-(\varepsilon I+\Sigma_{x})^{-1}\Bigr)E_{x}^{\top}.$ (24)
The diagonal entries of the matrix inside the parentheses are given by
$\varepsilon^{-1}-\frac{1}{\varepsilon+\sigma_{i}}=\frac{\sigma_{i}}{\varepsilon(\varepsilon+\sigma_{i})}\leq\frac{\sigma_{i}}{\varepsilon^{2}}.$ (25)
Therefore, eigencomponents corresponding to small eigenvalues $\sigma_{i}$ contribute negligibly to $R_{t}$. Moreover, approximating the empirical covariance $\sum_{s=1}^{t}x_{s}x_{s}^{\top}$ with rank $r$ before inversion $A_{t}$ or approximating only the correction $R_{t}$ with rank $r$ in decomposition for $A_{t}^{-1}$ leads to the same approximation of $A_{t}^{-1}$. This observation supports modeling $R_{t}$ using a low-rank representation when the empirical covariance matrix $\sum_{s=1}^{t}x_{s}x_{s}^{\top}$ has low effective rank, a property commonly employed in the literature on sketching methods kuzborskij2019efficient, chen2020efficient. Below we describe how to maintain a low-rank approximation of $R_{t}$ by controlling the ranks of $U_{t}$ and $V_{t}$ using the Projector-Splitting Integrator lubich2014projector. The motivation for this particular approximation method is discussed in the next section.
@@ 3 Projector-Splitting Integrator
We now aim to improve LinUCB using the representation in (21)–(2). In this form, the matrices $U_{t}$ and $V_{t}$ grow with time. At the same time, the key element of the representation (21) is the time-dependent matrix $D_{t}=U_{t}V_{t}^{\top}\in\mathbb{R}^{d\times d}$, whose evolution defines the dynamics of the representation. A natural approach is to approximate $D_{t}$ via a rank-$r$ truncated SVD with,
@@
At the same time, the orthogonal factors $(\bar{U}_{t},\bar{V}_{t})$ and $(\bar{U}_{t+1},\bar{V}_{t+1})$, corresponding to the matrices $D_{t}$ and $D_{t+1}$, are not guaranteed to be close, which might yield additional computational instability. Instead, we propose to rely on the projector-splitting integrator (PSI) approach of lubich2014projector, which constructs a dynamical low-rank approximation by solving
@@
where $\dot{D}_{t}=\frac{\mathrm{d}}{\mathrm{d}t}D_{t}$. The PSI method enables efficient iterative updates without explicitly forming $\tilde{D}_{t}$. The approximation is maintained in factorized form,
@@
where the factors $\tilde{U}_{t},\tilde{V}_{t}\in\mathbb{R}^{d\times r}$ have orthonormal columns, and $S_{t}\in\mathbb{R}^{r\times r}$ is invertible. Then we obtain factors $(\tilde{U}_{t+1},S_{t+1},\tilde{V}_{t+1}^{\top})$, such that
@@
We provide implementation details in Algorithm˜3.
We integrate this procedure into LinUCB with batch updates as follows. During training, the factors $U_{t}$ and $V_{t}$ are incrementally expanded using (2) or (44) until their number of columns exceed a predefined threshold $r$. At this time $t_{0}$, we compute a rank-$r$ SVD of $U_{t_{0}}V_{t_{0}}^{\top}$ to initialize the PSI factors. Subsequent updates apply Algorithm˜3 to maintain a fixed-rank approximation. The update increments $\Delta D_{t+1}$ are obtained directly from the LinUCB update rules. For rank-one updates,
@@
$\Delta D_{t+1}=Q_{t+1}(M_{t+1}-I)M_{t+1}^{-1}Q_{t+1}^{\top}(I-U_{t}V_{t}^{\top}).$ (31)
The complete integration of PSI into LinUCB is summarized in Algorithm˜1. We highlight that we do not need to form the matrices $L_{t}^{-1}$ given in (21) explicitly.
Complexity analysis. We analyze the computational complexity of the proposed PSI-LinUCB algorithm during training. The complexity of the PSI update
@@
does not depend on the number of interactions $B=B_{t,a}$ (that is, interactions with arm $a$ inside batch number $t$) and scales as $\mathcal{O}(dr^{2})$. This factor comes from the QR decomposition applied to $\mathbb{R}^{d\times r}$ matrices. When the rank exceeds the threshold, an SVD is computed once per arm (line 11 in Algorithm˜2) with complexity $\mathcal{O}(d(r+B)^{2}+(r+B)^{3})$, obtained via QR decompositions of factors of size at most $\mathbb{R}^{d\times(r+B)}$. Additionally, line 5 in Algorithm˜2 incurs a cost of $\mathcal{O}(dB^{2}+B^{3})$ due to QR and Cholesky decompositions. Thus, the overall complexity of handling blocks in the Update_arm algorithm (Algorithm˜2) is
@@
$\begin{split}a_{t}=\arg\max_{a\in\mathcal{A}_{t}}\left(\theta_{a}^{\top}x_{t,a}+\alpha\|(I-U_{t}V_{t}^{\top})L_{0}^{-1}x_{t,a}\|\right)\;.\end{split}$
@@
Table 1: Computational complexity comparison across algorithms
Algorithm Time cost per round Space
LinUCB $O(d^{2})$ $O(d^{2})$
CBRAP $O(dm+m^{3})$ $O(dm)$
CBSCFD $O(dm)$ $O(dm)$
DBSLinUCB $O(dl_{B_{t}})$ $O(dl_{B_{t}})$
PSI-LinUCB $O(dr)$ $O(dr)$
@@
The primary computational cost of our algorithm arises during the training phase, which requires on average $\mathcal{O}(d(r^{2}/B+B)+r^{3}/B+B^{2})$ operations per interaction. By choosing $B\propto r$ and assuming $r\ll d$, this reduces to average complexity of $\mathcal{O}(dr)$ operations per iteration. To our knowledge, the $\mathcal{O}(dr)$ average complexity is the best computational cost achieved by LinUCB-based algorithms. This is comparable with the most sample-efficient implementations of existing algorithms chen2020efficient, wen2024matrix, summarized in Table 1.
Theoretical properties. Assume that the matrix $U_{t}V_{t}^{\top}$ from the exact LinUCB update representation Theorem˜2 has rank not exceeding $r$ for any $t\in\{0,\ldots,T\}$. Then theoretical guarantees for PSI integrator applied with the same rank $r$ lubich2014projector[Theorem 4.1] ensures exact integration, that is, Algorithm˜3 maintains the exact $L_{t}^{-1}$ and and hence the exact inverse $A_{t}^{-1}=L_{t}^{-\top}L_{t}^{-1}$. Consequently, the estimator $\hat{\theta}_{t}=A_{t}^{-1}b_{t}$ and the confidence bonus in (34) exactly matches those of the standard LinUCB algorithm (6). In this setting, PSI-LinUCB admits the standard regret scaling of LinUCB under linear bandit assumptions of order $\tilde{O}(d\sqrt{T})$ with high probability after $T$ iterates, see abbasi2011improved.
Investigating the setting when the exact LinUCB algorithm matrix $U_{t}V_{t}^{\top}$ has larger rank is an important direction for future work. Existing regret analysis of LinUCB and its sketching variants kuzborskij2019efficient rely on the monotonicity properties of the estimates of feature covariance matrix. In our setting, we directly approximate the inverse regularized design matrix, and such monotonicity does not hold, making standard techniques unavailable.
@@
5:   $X_{t,a}.append\,[x_{t,a}]$, $R_{t,a}.append\,[r]$
6:  end for
7:  for each arm $a$ in $batch_{t}$ do
8:   $U_{t+1,a},V_{t+1,a},\theta_{t+1,a}\leftarrow$ Update_arm($a$)
9:  end for
@@
11:return $\theta_{t,a}$, $U_{t,a},V_{t,a}$ for each arm $a\in\mathcal{A}$.
@@ 4 Experimental Setup
We evaluate the proposed PSI-LinUCB against five baselines covering exact updates, sketching and random-projection approaches:
  • LinUCB li2010contextual: the standard LinUCB algorithm with batched updates;
  • LinUCB Classic  Angioli2025Efficient: LinUCB with exact rank-1 inverse updates via the Sherman–Morrison formula;
  • CBSCFD  chen2020efficient: a sketching-based method that approximates the feature covariance (design) matrix.
  • CBRAP yu2017cbrap: method based on random projections applied to feature covariance (design) matrix.
  • DBSLinUCB wen2024matrix: an adaptive sketching method that dynamically adjusts the sketch size over time.
We provide more details about existing methods in the Appendix, Appendix˜B.
@@ 4.1 Training Protocol
In order to run our algorithm on the datasets, we convert them into the bandit style environment following the pipeline below:
@@
In the sections below, we consider the datasets described in Table˜2. The LinUCB algorithm is implemented with batch processing to ensure scalability with respect to high-dimensional contextual features. User-item features are extracted via SVD decomposition of the user-item interaction matrix, with rank $r^{\prime}$ selected through spectral analysis during warm-up. To obtain high-dimensional contexts, the context vectors are constructed as outer products of user and item embeddings. All hyperparameters are tuned via cross-validation for each dataset and model, see Section˜C.1 in Appendix.
5 Scalability results
Increasing context size. To ensure the scalability of PSI-LinUCB, we vary the dimensions of context vectors $x_{t,a}$. We infer user-item features from SVD with varying rank $r^{\prime}$ on the subset of the Amazon Health dataset, and measure the computational time and memory usage. As shown in Figure˜1, LinUCB and LinUCB Classic fail to scale beyond moderate context sizes due to memory and time requirements for storing and inverting the full matrix $A_{t,a}$. In contrast, PSI-LinUCB operates efficiently even for large $d$ by maintaining only low-rank factors with $\mathcal{O}(dr)$ complexity. As shown in Table˜3, PSI-LinUCB achieves identical quality to LinUCB where both are feasible, while continuing to operate at larger context sizes. Warm-up phase results are provided in Appendix C, see Figure˜9.